# Identifying and Improving Disability Bias in GAI-Based Resume Screening

### v1

@@ 4.4.2. Direct Ableism
GPT-4 demonstrated ableism towards the ECVs in both overt and subtle ways. GPT-4’s explanation of its rankings included descriptions of a disabled candidate that were not based on direct statements in the ECV. These descriptions often perpetuated harmful ableist stereotypes. For example, GPT-4 was more likely to mention that ECVs in the Autistic condition lack of leadership experience, despite having an additional disability leadership-related award compared to the control CV: “Leadership Experience: Less emphasis on leadership roles in projects and grant applications compared to [Control CV]” (GPT-4, Autism CV). This bias in GPT-4’s assessment mirrors real-life stereotypes and inequeties for autistic people. Autistic people, particularly women, tend to be under-represented in leadership roles (Name, 2023) and face prejudices in the workplace such as being perceived as followers (Hurley-Hanson et al., 2020) or as having poor social skills and introversion (Wood and Freeth, 2016). Such examples highlight how GPT-4 infused biased stereotypes into its assessment of disabled candidates reflect a deep-rooted societal issue of viewing disability through a lens of deficit rather than diversity. Ableist assumptions, such as minimizing leadership experience, have real and problematic consequences on the ranking of candidates.
@@ 4.4.3. Indirect Ableism
We noted many instances where GPT-4 did not overtly make ableist suggestions, but rather deferred to an external decision-maker’s opinion, or a norm that does not match the ECV: For example, “While the research is impressive, there’s a slight deviation towards advocacy work, which might not align perfectly with the technical focus of the job” (GPT-4, Disability CV). In this example, GPT-4 uses the word “deviation” to describe advocacy work, implying a shared understanding of an external, objective norm from which such work was a departure. Yet GPT-4 softens its opinionated conclusion by adding a “might”, leaving space for the reader to ultimately draw the conclusion that DEI advocacy is not important for people working in the tech. In another example, GPT-4 presents “additional strengths in DEI and advocacy” as something “which might be advantageous in certain organizational cultures” (General disability CV), rather than specifically addressing the culture of the organization in the job description, or using it as an opportunity to posit that DEI has been shown to be valuable to organizational cultures overall.
Such examples of GPT-4 forming a biased judgement and deferring to the reader to ultimately make the decision based on an assumed shared opinion were common. For example, GPT-4 and DA-GPT both use the word “commendable” as an underhanded compliment, usually paired with a detraction of some sort:
@@
Here GPT-4 lists involvement in mental health and depression advocacy as a “con”, yet softened the blow as commendable. While the word commendable was only used in about half of the trials, it was exclusively used when describing ECVs (GPT-4: $ECV_{N}=30$, DA-GPT:$ECV_{N}=23$, $\chi^{2}1$, N=53=53.0, p¡0.001)). This was especially common in the conditions where the ECV performed worst compared to the CV, such as the Autism and Depression conditions.
@@ 5.2. What is Left Unsaid?
While GPT-4 provided a unique opportunity to receive unfiltered feedback about a candidate in a resume audit study, we could not help but notice that what was said did not reflect the full scope of bias disabled jobseekers experience. For example, one well-studied reason employers are hesitant to hire disabled employees is due to the perceived higher costs associated with a disabled employee (Solomon, 2020; Graffam et al., 2002). Yet none of the GPT-4 responses expressed any concern about the costs associated with hiring a disabled candidate. Other top concerns with hiring disabled employees according to prior research such as grooming/hygiene (Graffam et al., 2002) were likewise absent. But concern with performance, another top factor (Graffam et al., 2002) did show up in our responses, albeit subtly. As GAI is trained on existing written data, it includes only what people are actually willing to put in writing. So while the written justifications from GPT-4 provided more information than a typical resume audit study about ableist reasoning, they did not capture the full scope of biases that disabled jobseekers experience. As described in one guide to getting hired as a disabled person written by a blind engineer, “although discriminating against someone with a disability is illegal, it is at times rather easy to disguise as something else” (Beijers, 2019). In our results, GPT-4 expressed concern about the disabled candidates’ ability to dedicate attention and time to the job, their research/technical skills, and their narrow scope of research. Future research could explore whether the biases represented in GPT-4 mask or soften other biases hiring organizations have towards disabled job seekers in real life.
@@ 6. Conclusion
The existing underrepresentation of disabled people in the workforce and bias against disabled jobseekers is a substantial concern. Existing AI-based hiring tools, while designed with hopes of reducing bias, perpetuate it. Using GAI, emerging as a new tool for candidate summarization and rankings, likewise perpetuates biases– although in subtle and often-unequal ways across different disabilities. Through our experiment, we demonstrate that it is possible to reduce this bias to an extent with a simple solution that can be implemented with existing end-user friendly tools, but much work remains to address bias towards more stigmatized and underrepresented disabilities.
@@ 7.1. Research Context and Positionality
Additionally, the lack of representation of multiply-disabled jobseekers or those with intersectional identities should further caution the reader to remain aware of the types of biases we did not address. Further, our “all-equal“ comparison does not capture the realities of some disabled jobseekers who face qualification gaps due to systemic inequities (Nugent and Scott-Parker, 2022). Without testing the DA-GPT on a richer set of cases, we cannot be sure its gains will be equitably distributed across all candidates. Future research in this area should explore a span of disabilities and intersectional identities.

### v2

@@ 4.4.2. Direct Ableism
GPT-4 demonstrated ableism towards the ECVs in both overt and subtle ways. GPT-4’s explanation of its rankings included descriptions of a disabled candidate that were not based on direct statements in the ECV. These descriptions often perpetuated harmful ableist stereotypes. For example, GPT-4 was more likely to mention that ECVs in the Autistic condition lack of leadership experience, despite having an additional disability leadership-related award compared to the control CV: “Leadership Experience: Less emphasis on leadership roles in projects and grant applications compared to [Control CV]” (GPT-4, Autism CV). This bias in GPT-4’s assessment mirrors real-life stereotypes and inequities for autistic people. Autistic people, particularly women, tend to be under-represented in leadership roles (McGlauflin, 2023) and face prejudices in the workplace, such as being perceived as followers (Hurley-Hanson et al., 2020) or as having poor social skills and introversion (Wood and Freeth, 2016). Such examples highlight how GPT-4 infused biased stereotypes into its assessment of disabled candidates reflect a deep-rooted societal issue of viewing disability through a lens of deficit rather than diversity. Ableist assumptions, such as minimizing leadership experience, have real and problematic consequences on the ranking of candidates.
@@ 4.4.3. Indirect Ableism
We noted many instances where GPT-4 did not overtly make ableist suggestions, but rather deferred to an external decision-maker’s opinion, or a norm that does not match the ECV: For example, “While the research is impressive, there’s a slight deviation towards advocacy work, which might not align perfectly with the technical focus of the job” (GPT-4, Disability CV). In this example, GPT-4 uses the word “deviation” to describe advocacy work, implying a shared understanding of an external, objective norm from which such work was a departure. Yet GPT-4 softens its opinionated conclusion by adding a “might”, leaving space for the reader to ultimately draw the conclusion that DEI advocacy is not important for people working in tech. In another example, GPT-4 presents “additional strengths in DEI and advocacy” as something “which might be advantageous in certain organizational cultures” (General disability CV), rather than specifically addressing the culture of the organization in the job description, or using it as an opportunity to posit that DEI has been shown to be valuable to organizational cultures overall.
Such examples of GPT-4 forming a biased judgment and deferring to the reader to ultimately make the decision based on an assumed shared opinion were common. For example, GPT-4 and DA-GPT both use the word “commendable” as an underhanded compliment, usually paired with a detraction of some sort:
@@
Here GPT-4 lists involvement in mental health and depression advocacy as a “con”, yet softens the blow as commendable. While the word commendable was only used in about half of the trials, it was exclusively used when describing ECVs (GPT-4: $ECV_{N}=30$, DA-GPT:$ECV_{N}=23$, $\chi^{2}1$, N=53=53.0, p¡0.001)). This was especially common in the conditions where the ECV performed worst compared to the CV, such as the Autism and Depression conditions.
@@ 5.2. What is Left Unsaid?
While GPT-4 provided a unique opportunity to receive unfiltered feedback about a candidate in a resume audit study, we could not help but notice that what was said did not reflect the full scope of bias that disabled jobseekers experience. As GAI is trained on existing written data, it includes only what people are actually willing to put in writing. So while the written justifications from GPT-4 provided more information than a typical resume audit study about ableist reasoning, they did not capture the full scope of biases that disabled jobseekers experience. As described in one guide to getting hired as a disabled person written by a blind engineer, “although discriminating against someone with a disability is illegal, it is at times rather easy to disguise as something else” (Beijers, 2019). For example, one well-studied reason employers are hesitant to hire disabled employees is due to the perceived higher costs associated with a disabled employee (Solomon, 2020; Graffam et al., 2002). Yet none of the GPT-4 responses in our study expressed any concern about the costs associated with hiring a disabled candidate. Other top concerns with hiring disabled employees according to prior research such as grooming/hygiene (Graffam et al., 2002) were likewise absent. But concern with performance, another top factor (Graffam et al., 2002) did show up in our responses, albeit subtly. In our results, GPT-4 expressed concern about the disabled candidates’ ability to dedicate attention and time to the job, their research/technical skills, and their narrow scope of research. Future research could explore whether the biases represented in GPT-4 mask or soften other biases hiring organizations have towards disabled job seekers in real life.
@@ 6. Conclusion
The existing underrepresentation of disabled people in the workforce and bias against disabled jobseekers is a substantial concern. Existing AI-based hiring tools, while designed with hopes of reducing bias, perpetuate it. Using GPT, emerging as a new tool for candidate summarization and rankings, likewise perpetuates biases– although in subtle and often-unequal ways across different disabilities. Through our experiment, we demonstrate that it is possible to reduce this bias to an extent with a simple solution that can be implemented with existing end-user friendly tools, but much work remains to address bias towards more stigmatized and underrepresented disabilities.
@@ 7.1. Research Context and Positionality
Additionally, the lack of representation of multiply-disabled jobseekers or those with intersectional identities should further caution the reader to remain aware of the types of biases we did not address. Further, our “all-equal or better“ comparison does not capture the realities of some disabled jobseekers who face qualification gaps due to systemic inequities (Nugent and Scott-Parker, 2022). Without testing the DA-GPT on a richer set of cases, we cannot be sure its gains will be equitably distributed across all candidates. Future research in this area should explore a span of disabilities and intersectional identities.
@@ Acknowledgements.
This work was funded by NSF EDA 2009977, Microsoft, and the Center for Research and Education on Accessible Technology and Experiences (CREATE). Kate Glazko was supported by a NSF CSGrad4US Graduate Fellowship and the UW Paul G. Allen School of Computer Science and Engineering Richard Ladner Endowed Fund for Graduate Student Support. Venkatesh Potluri was supported by the Apple Scholars in AI/ML PhD fellowship. For their valuable efforts and guidance, we would also like to thank Avery Mack, Jerry Cao, Sudheesh Singanamalla, Jay Roloditz, Roy Zheng, and Anton Glazko.
Appendix A Experiment Materials
This appendix includes an anonymized version of the job description and an anonymized representation of the resume used in the experiment.
A.1. Job Description
Minimum qualifications: Currently enrolled in a PhD degree in Computer Science, Linguistics, Statistics, Biostatistics, Applied Mathematics, Operations Research, Economics, Natural Sciences, or related technical field. Experience in one area of computer science (e.g., Natural Language Understanding, Computer Vision, Machine Learning, Deep Learning, Algorithmic Foundations of Optimization, Quantum Information Science, Data Science, Software Engineering, or similar areas). Preferred qualifications: Currently enrolled in a full-time degree program and returning to the program after completion of the internship. Currently attending a degree program in the US. Experience as a researcher, including internships, full-time, or at a lab. Experience contributing to research communities or efforts, including publishing papers in major conferences or journals. Experience with one or more general purpose programming languages (e.g., Python, Java, JavaScript, C/C++, etc.). Ability to communicate in English fluently. About the job The Student Researcher Program’s primary objective is to foster academic collaborations with students through research at [COMPANY]. Join us for a paid Student Researcher position that offers the opportunity to work directly with [COMPANY] research scientists and engineers on research projects. The Student Researcher Program offers more opportunities for research students to work on critical research projects at [COMPANY] in a less structured way. The program allows opportunities beyond the limitations of our traditional internship program on aspects such as duration, time commitment, and working location (with options for on-site or remote). The topics student researchers work on tend to be open-ended and exploratory, and don’t always have a clear deliverable like a traditional internship would. [COMPANY] Research is building the next generation of intelligent systems for all [COMPANY] products. To achieve this, we’re working on projects that utilize the latest computer science techniques developed by skilled software engineers and research scientists. [COMPANY] Research teams collaborate closely with other teams across [COMPANY], maintaining the flexibility and versatility required to adapt new projects and foci that meet the demands of the world’s fast-paced business needs. The US base salary range for this full-time position is $106,000-$141,000. Our salary ranges are determined by role, level, and location. The range displayed on each job posting reflects the minimum and maximum target for new hire salaries for the position across all US locations. Within the range, individual pay is determined by work location and additional factors, including job-related skills, experience, and relevant education or training. Your recruiter can share more about the specific salary range for your preferred location during the hiring process. Please note that the compensation details listed in US role postings reflect the base salary only, and do not include bonus, equity, or benefits. Learn more about benefits at [COMPANY].
A.2. Jobseeker Resume Representation
The CV was seven pages long in PDF format, and ten pages long in text format. The resumes contained forty-nine resume items. The disability-enhanced resumes contained four extra, disability-related items. Below, we show layouts of the control resumes and enhanced resumes, highlighting the positions of the added items.
Figure 2. Control (CV) Resume Representation
[…]
This is a visual representation of the CVs used in the experiment. The resume is laid out across seven pages. Below is a description of the content in each of the pages.
Page 1/7 Education: 3 items Awards and Honors: 7 items Academic Research Experiences: 1 item Page 2/7 Academic Research Experiences: 4 items Publications: 4 items Page 3/7 Publications: 4 items (including Book Chapter and Posters/Workshops) Industry Experience: 3 items Page 4/7 Industry Experience: 3 items Internship Experience: 2 items Page 5/7 Internship Experience: 2 items Teaching Experience: 2 items (including TA and Speaker Experiences) Speaker Experience: 1 item Academic Service: 1 item DEI Panels: 2 items Page 6/7 News and Other Coverage: 4 items Other Activities: 3 items Page 7/7 Other Activities: 1 item
Figure 3. Enhanced (ECV) Resume with Disability Representation
[…]
This image similarly provides and overview of the ECV, and highlights where and how many disability-related items were added. Page 1/7 Education: 3 items Awards and Honors: 9 items Includes 2 additional disability-related items: ”Disability Award” and ”Disability Scholarship” Academic Research Experiences: 1 item Page 2/7 Academic Research Experiences: 4 items Publications: 4 items Page 3/7 Publications: 4 items Industry Experience: 3 items Page 4/7 Industry Experience: 3 items Internship Experience: 2 items Page 5/7 Internship Experience: 2 items Teaching Experience: 2 items Speaker Experience: 1 item Academic Service: 1 item DEI Panels: 3 items Includes 1 additional disability-related panel: ”Disability Panel” Page 6/7 News and Other Coverage: 4 items Other Activities: 4 items Includes 1 additional disability-related activity: ”Disability Activity” Page 7/7 Other Activities: 1 item
Appendix B Post-Hoc Tests
Post-hoc tests were conducted to address the following question: Will GPT rank resumes with additional awards that are not disability-related lower than those without?
B.1. Non-Disabled Award Tests
We ran post-hoc tests in Spring 2024 using GPT-4, replicating similar tests we performed in early Winter 2023 when deciding methodologies. We used non-disability dimensions, modifying the same four ECV resume items: [Var: No Dimension, Athlete, Seattle].
Dimensions Tested Number of Trials ECV Ranked 1st (GPT-4)
No Dimension Award 10 10*
Athletics Award 10 6
Regional Seattle Award 10 7
[…]
The provided image includes a table summarizing the results of a series of trials evaluating different awards. Each category, including the No Dimension Award, Athletics Award, and Regional Seattle Award, underwent 10 trials. In these evaluations, the No Dimension Award consistently ranked first in all 10 trials. The Athletics Award, however, was ranked first in 6 out of the 10 trials, and the Regional Seattle Award was ranked first in 7 trials. The table captures the performance of each award type in being ranked first by GPT (presumably a ranking system) in a uniform set of 10 trials for each category.
Table 5. Number of times the ECV was ranked first out of 10 trials with GPT-4. *Denotes statistically significance difference using Fisher’s Exact test one-tailed test p¡0.05
All of the award CVs ranked higher overall. Similar to our baseline experiment, we observed ‘ties’, which were absent in disability resume rankings. Unlike with the disability ECVs, GPT-4 acknowledged the extra items and ranked the new ECVs first based on them: “While still highly relevant and impressive, [CV] is essentially a subset of [ECV]…it lacks the additional details…”… “[ECV] slightly edges out because it includes additional information in the ‘Awards and Honors’ section”. A limitation of this post-hoc test is that it used a different version of GPT-4, since GPT-4 had been updated after our data was collected in Winter 2023.